# Systems-Level Attack Surface of Edge Agent Deployments on IoT

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2026conference: 6th Workshop on Machine Learning and Systems (EuroMLSys), co-located with EuroSys ’26; May 2026; Edinburgh, United Kingdomisbn: 978-1-4503-XXXX-X/2026/05ccs: Security and privacy Distributed systems securityccs: Computer systems organization Embedded and cyber-physical systemsccs: Security and privacy Network security
@@ 1. Introduction
LLM-based agents are rapidly moving from cloud APIs to edge hardware controlling physical devices. Commercial deployments now include SwitchBot AI Hub (local agent for smart home, launching Feb 2026) (SwitchBot, 2026), Home Assistant with MCP integrations enabling agentic automation (Home Assistant Community, 2025), and distributed frameworks like OpenClaw running on consumer hardware (OpenClaw Project, 2026). This shift changes the threat model fundamentally: where cloud-hosted agents present a single point of compromise with inherent data exfiltration, edge deployments distribute both capability and attack surface across heterogeneous nodes (Roman et al., 2018).
Recent security incidents highlight the urgency. Over 40,000 exposed OpenClaw gateway instances were discovered in late 2025, with 1,184 malicious packages uploaded to ClawHub (the agent skill repository) (OpenClaw, 2025). The 7,000-unit DJI Romo vacuum breach (Feb 2026) demonstrated how compromised credentials grant mass control over physically distributed agents (Hollister, 2026). Yet the systems community has not systematically examined how deployment architecture shapes agent security for IoT.
Critically, these agents do not choose their communication infrastructure: they inherit it. MQTT pub/sub is the de facto coordination layer in IoT deployments (AWS IoT Core, Azure IoT Hub, Home Assistant, SwitchBot), handling billions of device messages daily (Andy et al., 2017; Harsha et al., 2018). When LLM agents are deployed on this hardware, MQTT becomes their command-and-control plane by default. Our analysis examines the security consequences of this inheritance: what happens when a protocol designed for telemetry becomes the coordination bus for autonomous agents with physical actuation.
Existing agent security research focuses on prompt injection (Greshake et al., 2023; Debenedetti et al., 2024), tool misuse (Ruan et al., 2023), and model alignment (Shinn et al., 2023; Yao et al., 2022), assuming centralized cloud orchestration (Zhan et al., 2024). These defenses do not transfer cleanly to edge deployments where: (i) agents run on resource-constrained hardware unable to execute full guardrail stacks locally; (ii) network partitions isolate edge nodes, creating audit gaps; (iii) local data sovereignty eliminates cloud exfiltration risk but enables lateral movement between co-located agents; and (iv) physical actuation latency becomes a security-relevant metric (Alrawi et al., 2019).
@@
    Empirical findings. Edge-local MQTT pub/sub achieves 23 ms mean actuation-to-audit delay (P95 << 27 ms), enables complete provenance for cooperative agents but provides zero cryptographic enforcement, eliminates external data egress entirely during normal operation but silently degrades when fallback mechanisms trigger, and exposes vulnerability windows during network failover.
@@ 2.1. Deployment Architectures
The agent runtime (inference, tool calls, decision loop) executes in cloud infrastructure; IoT devices are controlled via cloud-to-device APIs. All sensor data, commands, and reasoning context transit externally. The architecture provides centralized audit and a single enforcement point, but inherently exfiltrates operational data during WAN partition (Roman et al., 2018).
@@
$2.$ Testbed
We deploy an actively-used edge-local agent swarm with different LLM models utilizing Openclaw  (OpenClaw, 2025) to control real IoT hardware (operational daily for home automation and research coordination):
@@ 3.1. Actuation-to-Audit Delay
MQTT round-trip latency: Rupert publishes $\to$ Percy receives and echoes $\to$ Rupert records. Three payload sizes (50 B, 1 KB, 10 KB), N=150N{=}150 messages per class, microsecond-precision timestamps.
Mac mini $\to$ Moto G35 (via Tailscale, N=150N{=}150 per size).
Table 2. Actuation-to-audit delay: Mac mini $\to$ Moto G35 via Tailscale (N=150N{=}150 per payload size).
@@
50 B 23.6 ms 23.4 ms 26.9 ms 33.0 ms
1 KB 22.5 ms 22.4 ms 26.6 ms 30.6 ms
10 KB 34.5 ms 34.2 ms 41.9 ms 44.6 ms
NUC $\to$ Mac mini (cross-validation, N=50N{=}50 per size).
Table 3. Cross-validation latency: NUC $\to$ Mac mini (N=50N{=}50 per payload size).
[…]
128 B 64.4 ms 65.0 ms
1 KB 64.3 ms 64.8 ms
8 KB 64.5 ms 65.1 ms
The $2.7{\times}$ latency difference between paths (Tables 2 and 3) reflects node heterogeneity: the NUC’s low-power N150 processor (single performance core) versus the Mac mini’s M4 (8 cores, unified memory architecture). Notably, the NUC path shows near-zero variance across payload sizes (64.3–64.5 ms), suggesting a fixed scheduling overhead rather than serialization cost. Both paths remain well within the interception window for physical actuators.
@@
Under sustained load on the NUC (N=100N{=}100 messages at 128 B, rapid-fire), mean latency is 49.8 ms with 0% degradation. It confirms that the MQTT broker handles burst traffic without queuing delays even from the most constrained node.
@@
Real-time safety monitors are feasible for edge-local MQTT deployments. A monitor subscribing to the audit topic observes every action within ${\sim}23$ ms on the fastest path (64 ms on the NUC path). It is within the actuation window of electromechanical devices such as door locks (${\sim}500$ ms) and motorized valves (${\sim}1$22 s), though insufficient for solid-state relays (<5{<}5 ms).
@@ 3.2. Provenance Chain Completeness
MQTT provides transport-level delivery but no application-level security enforcement (Table 4). Provenance therefore depends entirely on agent cooperation. Message signing (HMAC/JWT), nonce-based replay protection, and topic access-control lists (ACLs) are required yet absent from the default deployment. None of the open-source agent frameworks we surveyed implement them either.
$3.$ Data Sovereignty and Egress
@@
Table 5. Data egress comparison: edge-local vs. cloud architecture during a 10-minute operational session.
@@
Edge-local deployments exhibit zero external data egress during normal operation (Table 5). All communication remains within the Tailscale mesh. The operation counts differ because our cloud baseline captures only inference API calls, not a full cloud-orchestrated workload performing identical tasks. The comparison demonstrates the structural egress difference rather than a controlled workload-matched measurement. Even so, a minimal cloud workload produces 65 KB of egress, while more edge-local operations produce zero. This eliminates structural data exfiltration as an attack vector: an adversary cannot observe agent context by monitoring cloud API traffic but must compromise a node within the mesh.
@@ 4.1. Architecture IS Security Posture
  • An edge-local agent with no guardrails produces zero external egress and enables 23 ms actuation-to-audit delay, sufficient for real-time interception of most physical actuators.
@@ 4.2. The Provenance Gap
This gap is architectural, not incidental. MQTT was designed as a lightweight pub/sub transport, not a secure command-and-control protocol. Closing the gap requires: (i) per-agent signing keys with broker-side verification; (ii) nonce-based replay protection; and (iii) topic-level ACLs enforced at the broker. These are well-understood techniques, but their absence from current agent frameworks reflects a community assumption that agents operate in trusted environments, an assumption that IoT deployments invalidate (OWASP Foundation, 2025; Alrawi et al., 2019). Scaling beyond our three-agent testbed to NN agents will require formal trust policies: capability delegation and least-privilege topic access. None of these are provided by MQTT natively.
@@ 4.3. Emergent Coordination Failures
$4.$ Silent Sovereignty Degradation
@@ 5. Related Work
Greshake et al. (Greshake et al., 2023) and Zhan et al. (Zhan et al., 2024) demonstrated prompt injection and surveyed agent attacks, both assuming cloud-hosted architectures. Xi et al. (Xi et al., 2025) proposed guardrail taxonomies without addressing edge deployment constraints. AgentDojo (Debenedetti et al., 2024) and LM-emulated sandboxes (Ruan et al., 2023) evaluate agent vulnerabilities in controlled settings; OWASP (OWASP Foundation, 2025) and PASB (Wang et al., 2026) catalogue and benchmark agentic risks. These works focus on what agents do wrong (tool misuse, prompt compliance); we complement them by examining how the infrastructure agents inherit shapes the attack surface independent of model behaviour.
@@
Edge computing security surveys (Shi et al., 2016; Roman et al., 2018) identify data sovereignty as a central concern; we measure it empirically and show it degrades silently under resource exhaustion. IoT security work has focused on device-level vulnerabilities (Meneghello et al., 2019; Alrawi et al., 2019) and smart-home automation risks (Surbatovich et al., 2017), but not the LLM agent layer that now mediates between user intent and physical actuation. MQTT-specific analyses (Andy et al., 2017; Harsha et al., 2018; Firdous et al., 2017; Mishra and Kertesz, 2020) document authentication and access control gaps in telemetry contexts; we show these gaps become safety-critical when MQTT carries autonomous agent commands to physical actuators. Our testbed uses WireGuard (Donenfeld, 2017) via Tailscale, but our findings confirm that transport-layer encryption does not address application-layer coordination vulnerabilities (S1a–c).
@@ 6. Conclusion
Through empirical measurements on a three-node edge-local agent swarm, we showed that deployment architecture, not model or prompt design, is the primary determinant of security posture for agent-controlled IoT. Edge-local MQTT eliminates structural data exfiltration and enables actuation-to-audit delay, but provides zero cryptographic provenance enforcement, exposes failover blind spots, and silently degrades data sovereignty under stress. Two emergent failures (coordination-state divergence and induced trust erosion) further demonstrate that unauthenticated message buses are insufficient for multi-agent coordination. Closing these gaps requires per-agent signing keys, sovereignty-aware fallback policies, and network-level fast roaming; none are provided by current agent frameworks. The systems community should treat where agents run and how they communicate as first-class security design choices, not implementation details subordinate to model-level alignment.

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2026journalyear: 2026conference: The 6th Workshop on Machine Learning and Systems; April 27–30, 2026; Edinburgh, Scotland Ukbooktitle: The 6th Workshop on Machine Learning and Systems (EuroMLSys ’26), April 27–30, 2026, Edinburgh, Scotland Ukdoi: 10.1145/3805621.3807636isbn: 979-8-4007-2605-7/26/04ccs: Security and privacy Mobile and wireless securityccs: Security and privacy Distributed systems securityccs: Computer systems organization Embedded and cyber-physical systems
@@ 1. Introduction
LLM-based agents are rapidly moving from cloud APIs to edge hardware controlling physical devices. Commercial deployments now include SwitchBot AI Hub (a smart-home hub advertising local AI capabilities) (SwitchBot, 2026), Home Assistant’s MCP integration exposing tools to conversation agents (Home Assistant, 2026), and distributed frameworks like OpenClaw running on consumer hardware (OpenClaw Project, 2026a). This shift changes the threat model fundamentally: where cloud-hosted agents present a single point of compromise with inherent data exfiltration, edge deployments distribute both capability and attack surface across heterogeneous nodes (Roman et al., 2018).
Recent deployments and incidents highlight the urgency. OpenClaw’s official security guidance explicitly centers prompt injection, gateway exposure, and trusted-operator assumptions as first-order risks in personal-assistant deployments (OpenClaw Project, 2026b). The 7,000-unit DJI Romo vacuum breach (Feb 2026) demonstrated how compromised credentials grant mass control over physically distributed agents (Hollister, 2026). Yet the systems community has not systematically examined how deployment architecture shapes agent security for IoT.
Critically, these agents do not choose their communication infrastructure: they inherit it. MQTT pub/sub is a widely used coordination substrate in IoT deployments (Mishra and Kertesz, 2020; Andy et al., 2017; Harsha et al., 2018). When LLM agents are deployed on this hardware, MQTT becomes their command-and-control plane by default. Our analysis examines the security consequences of this inheritance: what happens when a protocol designed for telemetry becomes the coordination bus for autonomous agents with physical actuation.
Existing agent security research focuses on prompt injection (Greshake et al., 2023; Debenedetti et al., 2024), tool-use risk evaluation (Ruan et al., 2023), and agent reasoning or self-improvement scaffolds (Shinn et al., 2023; Yao et al., 2022), typically in centralized digital environments rather than edge-local IoT deployments (Zhan et al., 2024). These defenses do not transfer cleanly to edge deployments where: (i) agents run on resource-constrained hardware unable to execute full guardrail stacks locally; (ii) network partitions isolate edge nodes, creating audit gaps; (iii) local data sovereignty eliminates cloud exfiltration risk but enables lateral movement between co-located agents; and (iv) physical actuation latency becomes a security-relevant metric.
@@
    Empirical findings. The full end-to-end actuation pipeline (MQTT publish through Home Assistant actuation to confirmation) achieves ${\sim}117$ ms mean delay (P95 <152{<}152 ms), enables complete provenance for cooperative agents but provides zero cryptographic enforcement, eliminates external data egress entirely during normal operation but silently degrades when fallback mechanisms trigger, and exposes vulnerability windows during network failover.
@@ 2.1. Deployment Architectures
The agent runtime (inference, tool calls, decision loop) executes in cloud infrastructure; IoT devices are controlled via cloud-to-device APIs. All sensor data, commands, and reasoning context transit externally. The architecture provides centralized audit and a single enforcement point, but necessarily places operational data on external infrastructure and expands confidentiality and trust dependencies (Roman et al., 2018).
@@
2.2. Testbed
We deploy an actively-used edge-local agent swarm with different LLM models utilizing OpenClaw (OpenClaw Project, 2026a) to control real IoT hardware (operational daily for home automation and research coordination):
@@ 3.1. Actuation-to-Audit Delay
We measure the full end-to-end actuation pipeline, not just the MQTT transport hop. An actuation command traverses four stages: (1) MQTT publish from the commanding agent to the broker, (2) Home Assistant state read (pre-actuation device check), (3) Home Assistant service call (physical actuation), and (4) Home Assistant confirmation (post-actuation state verification). We instrument each stage with microsecond-precision timestamps (N=150N{=}150).
Table 2. End-to-end actuation-to-audit delay: full pipeline from agent command to confirmed device state change (N=150N{=}150).
@@
MQTT publish 36.4 ms 36.0 ms 38 ms Agent $\to$ broker
HA state read 28.6 ms 29.0 ms 31 ms Pre-actuation check
HA actuation 27.6 ms 26.0 ms 64 ms Service call (light toggle)
HA confirmation 24.1 ms 24.0 ms 31 ms Post-actuation verify
Total E2E 116.7 ms 114.5 ms 152 ms Full pipeline
Pipeline decomposition.
The MQTT transport hop accounts for only 31% of the total delay (Table 2); the Home Assistant REST bridge and device confirmation contribute the remaining 69%. The HA actuation stage shows the highest variance (P95 of 64 ms vs. mean of 27.6 ms), reflecting device-dependent response times. The total E2E delay of ${\sim}117$ ms is $3.2{\times}$ the MQTT-only figure but remains well within the interception window for electromechanical actuators such as door locks (${\sim}500$ ms) and motorized valves (${\sim}1$22 s).
Cross-path validation.
The NUC $\to$ Mac mini MQTT path exhibits ${\sim}64$ ms mean latency (N=50N{=}50), a $2.7{\times}$ overhead reflecting node heterogeneity (NUC’s single-core N150 vs. Mac mini’s M4). The NUC path shows near-zero variance across payload sizes (64.3–64.5 ms), suggesting fixed scheduling overhead rather than serialization cost.
@@
Under sustained load on the NUC (N=100N{=}100 messages at 128 B, rapid-fire), mean latency is 49.8 ms with 0% degradation, confirming that the MQTT broker handles burst traffic without queuing delays even from the most constrained node.
@@
Real-time safety monitors are feasible for edge-local MQTT deployments. A monitor observes the full actuation pipeline within ${\sim}117$ ms (P95 <152{<}152 ms). This is within the actuation window of electromechanical devices such as door locks (${\sim}500$ ms) and motorized valves (${\sim}1$22 s), though insufficient for solid-state relays (<5{<}5 ms).
@@ 3.2. Provenance Chain Completeness
MQTT provides transport-level delivery but no application-level security enforcement (Table 3). These gaps are well-documented in telemetry contexts (Firdous et al., 2017; Harsha et al., 2018); the critical difference here is that MQTT now carries autonomous agent commands to physical actuators. A spoofed “unlock door” command is not a corrupted sensor reading: it triggers irreversible physical action within ${\sim}117$ ms (Table 2), faster than any human-in-the-loop can intervene. Provenance therefore depends entirely on agent cooperation. Message signing (HMAC/JWT), nonce-based replay protection, and topic ACLs are required yet absent from the default deployment. We did not observe these mechanisms in the open-source agent frameworks we used.
3.3. Data Sovereignty and Egress
@@
Table 4. Data egress comparison: edge-local vs. cloud architecture. The aggregate row reports a 10-minute session; the normalised row controls for operation count.
@@
[…]
Edge-local MQTT $\sim$}805 B 1
Cloud API $\sim$}3,244 B 1
Edge-local deployments exhibit zero external data egress during normal operation (Table 3.3). All communication remains within the Tailscale mesh. To control for the differing operation counts in our aggregate measurement, we also report normalised per-actuation egress for matched operation types: cloud API calls transmit $4.0{\times}$ the bytes of equivalent edge-local operations. The edge-local bytes represent intra-mesh MQTT traffic (never leaving the Tailscale network); the cloud bytes traverse the public internet to inference providers. This eliminates structural data exfiltration as an attack vector: an adversary cannot observe agent context by monitoring cloud API traffic but must compromise a node within the mesh.
@@ 4.1. Architecture IS Security Posture
  • An edge-local agent with no guardrails produces zero external egress and enables ${\sim}117$ ms end-to-end actuation-to-audit delay, sufficient for real-time interception of most physical actuators.
@@
4.2. Relation to Non-LLM Distributed Systems
Many of the vulnerabilities we identify are well-studied in robotics, microservices, and industrial SCADA (Supervisory Control and Data Acquisition) systems (Stouffer et al., 2011). They include unauthenticated messaging, state divergence, failover blind spots. Authenticated messaging (TLS client certificates, HMAC-signed payloads), shared-state protocols (distributed consensus), and network fast roaming (multi-path TCP) are mature techniques in those domains. We believe these transfer directly to the MQTT and network layers of LLM agent deployments. What does not transfer is the trust model: traditional distributed systems assume deterministic, programmed agents with fixed capabilities, whereas LLM agents exhibit non-deterministic behaviour, interpret natural-language instructions, and can be manipulated via prompt injection. The S1c vulnerability (induced trust erosion) is specific to this class of agent: a deterministic system would not “decide” to distrust its communication channel based on message content. Mitigations must therefore combine established distributed-systems techniques at the infrastructure layer with LLM-specific defences at the agent layer. Neither alone is sufficient.
@@ 4.3. The Provenance Gap
This gap is architectural, not incidental. MQTT was designed as a lightweight pub/sub transport, not a secure command-and-control protocol. Closing the gap requires: (i) per-agent signing keys with broker-side verification; (ii) nonce-based replay protection; and (iii) topic-level ACLs enforced at the broker. These are well-understood techniques, but their absence from the agent frameworks used in our deployment is consistent with security guidance that still assumes a trusted operator boundary or otherwise treats agentic security at a higher level than coordination-bus provenance (OpenClaw Project, 2026b; OWASP Foundation, 2025). Scaling beyond our three-agent testbed to $N$ agents will require formal trust policies: capability delegation and least-privilege topic access. None of these are provided by MQTT natively.
@@ 4.4. Emergent Coordination Failures
4.5. Silent Sovereignty Degradation
@@ 5. Related Work
Greshake et al. (Greshake et al., 2023) and Zhan et al. (Zhan et al., 2024) demonstrated indirect prompt injection against tool-integrated LLM applications and agents. Xi et al. (Xi et al., 2025) surveyed LLM-based agents and outlined their architectural components and open problems, but did not address edge deployment constraints. AgentDojo (Debenedetti et al., 2024) and LM-emulated sandboxes (Ruan et al., 2023) evaluate agent vulnerabilities in controlled settings; OWASP (OWASP Foundation, 2025) and PASB (Wang et al., 2026) catalogue and benchmark agentic risks. These works focus on what agents do wrong (tool misuse, prompt compliance); we complement them by examining how the infrastructure agents inherit shapes the attack surface independent of model behaviour.
@@
Edge computing and MEC surveys (Shi et al., 2016; Roman et al., 2018) discuss latency, locality, and security/privacy tradeoffs; we measure one such tradeoff empirically and show that sovereignty can degrade silently under resource exhaustion. IoT security work has focused on device-level vulnerabilities (Meneghello et al., 2019; Alrawi et al., 2019) and smart-home automation risks (Surbatovich et al., 2017), but not the LLM agent layer that now mediates between user intent and physical actuation. MQTT-specific analyses (Andy et al., 2017; Harsha et al., 2018; Firdous et al., 2017; Mishra and Kertesz, 2020) document authentication and access control gaps in telemetry contexts; we show these gaps become safety-critical when MQTT carries autonomous agent commands to physical actuators. Our testbed uses WireGuard (Donenfeld, 2017) via Tailscale, but our findings confirm that transport-layer encryption does not address application-layer coordination vulnerabilities (S1a–c).
@@ 6. Conclusion
Through empirical measurements on a three-node edge-local agent swarm, we showed that deployment architecture, not model or prompt design, is the primary determinant of security posture for agent-controlled IoT. Edge-local MQTT eliminates structural data exfiltration and enables low actuation-to-audit delay, but provides zero cryptographic provenance enforcement, exposes failover blind spots, and silently degrades data sovereignty under stress. Two emergent failures (coordination-state divergence and induced trust erosion) further demonstrate that unauthenticated message buses are insufficient for multi-agent coordination. Closing these gaps requires per-agent signing keys, sovereignty-aware fallback policies, and network-level fast roaming; we did not observe them in the frameworks used in our deployment or examined in this work. The systems community should treat where agents run and how they communicate as first-class security design choices, not implementation details subordinate to model-level alignment.
Acknowledgements.
This work was supported by the CHIST-ERA grant CHIST-ERA-22-SPiDDS-02 (GRAPHS4SEC). This work was conducted within the Networks and Systems Lab at Imperial College London. All adversarial experiments were conducted on the authors’ own hardware in a closed local network with no connection to production systems or third-party infrastructure.
Appendix A Extended Testbed and Measurements
Following the measurements reported in the main text (three-node testbed), the deployment expanded organically to five nodes. This appendix documents the extended topology and additional measurements that address scaling behaviour.
The original three-node topology (Figure 1) remains the basis for all main-text measurements. The two additional nodes described below joined the deployment subsequently.
A.1. Extended Topology
Two nodes were added to the original three-node swarm:
  • Pixel 4 (Android 14, Google Pixel 4, Termux): mobile agent “Darcy” (Grok 4.2). Darcy controls proprietary IoT applications that expose no external API, using UI automation through ADB and accessibility services. Darcy also accesses on-device sensors (camera, ambient light, gyroscope) that are physically unavailable to hub-resident agents, which is a key reason a single centralized agent cannot replace the distributed architecture. Darcy’s role mirrors Percy’s, enabling comparison across different models and hardware.
  • Intel NUC11 (x86, 4 cores, 7.5 GB RAM, Ubuntu 24.04): librarian agent “Dewey” (observational only). Dewey maintains a Git repository as the canonical shared record of system coordination state. It subscribes to mirrored MQTT traffic and records externally relevant events with structured metadata (timestamp, sender, outcome). Events include task dispatches, device-state transitions, conflict resolutions, and recovery outcomes. Dewey does not issue commands and holds no device credentials; its role is purely observational. The repository provides three properties: (1) versioned shared state that agents can inspect and recover after restart; (2) attribution and history preserved in a diffable timeline; and (3) stale-state detection via base_commit comparison against the current HEAD.
Device adapters.
Devices are not agents. Lights, speakers, cameras, and sensors are thin, deterministic endpoints controlled by managers through protocol-specific adapters. This keeps the agent count small and avoids an “agent per device” design.
Dewey as operational response to S1b.
The git-backed shared state plane maintained by Dewey directly addresses the coordination-state divergence vulnerability (S1b) identified in the main text (§3.2). Rather than relying on MQTT messages as the sole coordination medium (which produces divergent local copies that drift silently) agents can reference canonical state in the repository. Conflict detection becomes explicit through Git’s merge semantics rather than invisible.
A.2. Scaling: Fan-out Latency
To address the question of how the MQTT coordination layer scales beyond three nodes, we measure fan-out latency as the number of target agents increases from 1 to 4 (the full swarm excluding the sender).
Table 7. Fan-out latency: sequential publish to $k$ targets vs. broadcast (N=50N{=}50 per configuration).
[…]
1 target (original) 35 ms
2 targets (fan-out) 41 ms
4 targets (full swarm) 55 ms
Broadcast (single pub) 35 ms
Sequential fan-out grows linearly at ${\sim}5$ ms per additional target (Table 7), reflecting per-publish overhead rather than broker contention. Broadcast (a single MQTT publish to a shared topic) remains constant regardless of subscriber count, as the broker handles fan-out internally. For safety-critical commands that require all agents to observe the same action simultaneously, broadcast is both faster and semantically stronger. The linear sequential cost is modest for our five-node swarm but would become a bottleneck at N>20N{>}20 agents if per-agent addressing is required.